Determine whether or not the following are continuous functions.

1. Ζ’(π‘₯)=5π‘₯+3 2. Ζ’(π‘₯)=βˆ’4π‘₯+2

3. Ζ’(π‘₯)=2π‘₯2 +π‘₯βˆ’3

4. Ζ’(π‘₯)={2π‘₯βˆ’3 iΖ’ π‘₯β‰₯2 βˆ’2π‘₯ + 2 iΖ’ π‘₯ < 2

5. Ζ’(π‘₯)={|π‘₯+2|iΖ’π‘₯Gβˆ’2 4 iΖ’ π‘₯ = βˆ’2


1
Expert's answer
2022-03-20T10:12:10-0400

1. "\u0192(\ud835\udc65)=5\ud835\udc65+3"

Consider the difference

"f(x+\Delta x)-f(x)=5(x+\Delta x)+3 - (5x+3)=5\Delta x\to 0" as "\Delta x\to0", for any "x\in\mathbb{R}". Therefore, this function is continuous for any "x\in\mathbb{R}".


2. "\u0192(\ud835\udc65)=\u22124\ud835\udc65+2"

Consider the difference

"f(x+\Delta x)-f(x)=-4(x+\Delta x)+2 - (−4𝑥+2)=-4\Delta x\to 0" as "\Delta x\to0", for any "x\in\mathbb{R}". Therefore, this function is continuous for any "x\in\mathbb{R}".


3. "\u0192(\ud835\udc65)=2\ud835\udc65^2 +\ud835\udc65\u22123"

Consider the difference

"f(x+\Delta x)-f(x)=2(x+\Delta x)^2 +(x+\Delta x)−3 - (2𝑥^2 +𝑥−3)"

"=4x\Delta x+2(\Delta x)^2\to 0" as "\Delta x\to0", for any "x\in\mathbb{R}".

Therefore, this function is continuous for any "x\in\mathbb{R}".


4. "ƒ(𝑥)=\left\{
\begin{matrix}
2𝑥−3 & \rm{if}\, 𝑥≥2 \\
−2𝑥 + 2 & \rm{if}\, 𝑥 < 2
\end{matrix}
\right."

For any "x<2" consider the difference

"f(x+\Delta x)-f(x)=-2(x+\Delta x)+2 - (−2𝑥 + 2)=-2\Delta x\to 0"

as "\Delta x\to0". Therefore, this function is continuous for any "x<2".

For any "x>2" consider the difference

"f(x+\Delta x)-f(x)=2(x+\Delta x)+2 - (2𝑥 + 2)=2\Delta x\to 0"

as "\Delta x\to0". Therefore, this function is continuous for any "x>2".

For "x=2" consider one-sided limits:

"\lim\limits_{\Delta\to+0}f(2+\Delta x)=\lim\limits_{\Delta\to+0}(2(2+\Delta x)-3)=1"

"\lim\limits_{\Delta\to-0}f(2+\Delta x)=\lim\limits_{\Delta\to+0}(-2(2+\Delta x)+2)=-2"

The one-sided limits are not equal, therefore, the function is not continuous at "x=2".


5. "ƒ(𝑥)=\left\{
\begin{matrix}
|𝑥+2| & \rm{if}\, 𝑥\ne -2 \\
4 & \rm{if}\, 𝑥 = -2
\end{matrix}\right."

For any "x>-2" consider the limit

"\lim\limits_{\Delta\to 0}f(x+\Delta x)=\lim\limits_{\Delta\to0}|x+\Delta x+2|"

"=\lim\limits_{\Delta\to0}(x+\Delta x+2)=x+2=f(x)"

Therefore, the function is continuous for any "x>-2".

For any "x<-2" consider the limit

"\lim\limits_{\Delta\to 0}f(x+\Delta x)=\lim\limits_{\Delta\to0}|x+\Delta x+2|"

"=\lim\limits_{\Delta\to0}(-x-\Delta x-2)=-x-2=f(x)"

Therefore, the function is continuous for any "x<-2".

For "x=-2" consider one-sided limits:

"\lim\limits_{\Delta\to+0}f(-2+\Delta x)=\lim\limits_{\Delta\to+0}|-2+\Delta x+2|=0"

"\lim\limits_{\Delta\to-0}f(-2+\Delta x)=\lim\limits_{\Delta\to-0}|-2+\Delta x+2|=0"

They are not equal to "f(-2)=4". Therefore, the function is not continuous at "x=-2".